USA report points to remaining issues

by Chris Bergin

Just how much work is required to ensure Discovery launches in the new July window became apparent in Thursday’s internal e-mail to United Space Alliance (USA) employees.

The full internal e-mail follows:

Chairperson: Mr. Bill Parsons
* MMT Sim Final is done. The debrief went extremely well. Got out of it what they wanted from it.
* Completed the TCDT at KSC.

Integrated Flow Status (USA/KSC) OV-103 (STS-114)
* Completed the TCDT and it went well
* ECO Sensor troubleshooting and plan to be complete Noon tomorrow to support a 4pm meeting
* Continue to refine and complete schedule options to support the goal to launch in July

OV-104 (STS-121 and STS-300)
* Power-on System testing is about 70% complete
* Still working the LG door as rigging continues and working the cycles
* RSB is a good news story. Turned to GREEN due to Lessons Learned from OV-103.

* ET-121–Finished closeouts on SRB stacking and continue to support all of Lockheed requirements in the VAB

KSC Logistics (USA/KSC)
* ECO Sensor troubleshooting–going into Vib Test today on the Point Sensor Box and plan to wrap up today and get into Post-vib analysis tomorrow and get data to Orbiter to support meetings tomorrow

KSC Deputy Program Manager (NASA/KSC)
* Ray J–The Airspace Exercise that was supposed to be today looks like will slip to July 6th or 8th

SSME Testing (Rocketdyne/SSC)
* Had a successful test on Monday for Green run on the 9th Pump needed for first flight. The next test is 5/24 and is planned as a celebration of the 30 year anniversary of 1st hotfire test at Stennis for SSME

SSME (Rocketdyne/KSC)
* On OV-103, completed the gimbal bushing inspection and was good. Will start 104 tomorrow to complete the rest of the fleet next week.

* Pump tested at Stennis on Monday, this was the 9th pump and we had the data review yesterday and that looked good. As this is an acceptable pump, this gives us a 9th pump. Over the last 10-12 months, Pratt-Whitney has re-cycled 12 LOX pumps. Over the past 8 months, SSC has completed 11 tests on A2 and 13 on A1 for 24 total tests. This gives 10 Controllers and 9 LOX pumps needed for first flight. Have not missed a single scheduled test date for any problem at Stennis. Thanks to Stennis for doing their part in getting ready to fly. Mr. Parsons agreed and “hats off” to the whole SSME team.

* ET121 at KSC, LO2 Diffuser changeout, tank is vented. Starting the Ogive cover plate removal and should be off by the end of the shift tonight. Tracking to be complete with diffuser changed out and cover plate back on and the nose cone re-installed by the 12th.
* Bellows heater mod, putting the intertank access kit in and ready to start the harness installation in the intertank and will do as soon as get access.
* LH2 Diffuser–working on getting assests. Removals in-work at MAF and the 1st part available by the 9th. Putting schedule together to make the 2nd one available about 6 or 7 days after that. This will support different options at KSC.
* ET119–reworking schedule and Critical Path flows through the Intertank and are putting additional resources on it and moving back to left to support the Option 2 schedule
* Bringing to the Noon Board today, a request to pull the diffuser out of ET117 and R&R LH2 diffusers out of 120 and 121.

* Finished the PRA activities at MAF yesterday. Came up with an agreement on the Monte Carlo inputs.

MSFC Integration (NASA/MSFC)
* Robert Champion will have a meeting at 9am central on Tanking Test Objectives to be prepared for 3pm meeting this afternoon. Using what was talked in the 3pm meeting yesterday as a baseline and have received some inputs already.

MSFC Deputy Program Manager (NASA/MSFC)
* On the Noon Board item–Diffuser is consistent with plan that has been discussed.
* May end up changing stacks and wants all elements to look to see if have any issues. Parsons commented that no decision has been made yet and wants anyone who may have an issue to make it known.
* Option 1A and Option 2…do not need to make a decision until Friday afternoon on the options.

Orbiter (USA/JSC)
* Testing on SN 120 Thruster Injector to see if can determine the effects of Okite. There are some that are happy with the explanation and and those with concerns did not see anything in test to mitigate concerns. Results could be called inconclusive. Will look at some more tests.
* At Southwest Research, we continue to shoot Tyvek covers at windows. To achieve speeds that aero folks say will impact windows and so far have seen no damage to windows. Need to finish total number of shots to get the statistical significance.

Orbiter Project (NASA/JSC)
* Thruster 120 tagup at 3pm today for additional results for the concerns for Okite
* Continuing to work the HUD alignment with the attachments and the 1/4 inch freeplay. Plan is to R&R the Power Drive Unit (PDU) and work may be done today. If problem is in the alignment interface with the Orbiter, then would have to be an OPF activity. Before doing that, would have to talk to crew and MOD as to whether we would have to do that or not.
* DCS 760 cameras–We had another failure on board Station. There were originally 3 old configuration cameras to be used for RPM. One of those failed and we are down to 2. We shipped up on Progress 17P, 2 of the latest configuration units. One of those failed so we have 3 on board Station today and do not feel this is sufficient. We are working with ISS to launch one on 18P in June of the new config and that will give us 4. We really want 2 prime and 1 hot backup for the RPM for STS-114. There is a concern that Progress environment is part of the contributor to the anomalies and we are taking some precautions and working on re-packing camera before shipment to Russia.
* In Denver, STA54 material was dispensed that was on shelf for about 104 days. All samples tested came within the requirement for the hardness. Looks good for the shelf life on material.
* An “I Forgot” relative to the Blankets and the Hydraulic Fluid—The KSC team has been busy with removing the C9 Coating while the vehicle is still at the pad. Process seems to be working very well. Blankets had a level of contamination from the hydraulic fluid. Seeing a little bit of fraying of the blankets during the removal process. There is nothing at this point that concerns anyone that will drive us to R&R. Saw a few Open Actions that need to finish relative to contamination. Right now looks like we will be fine with just removing the C9 Coating. Testing at Marshall on removing and re-applying the coating after hydraulic fluid contamination is
not an issue.

Flight Ops (USA/JSC)
* Working informally on potential of the stack change. There will be some impacts on Flt Ops side. Assessments are not complete but we have different SRB burn rates and mass properties which influence a lot of the things we deliver from Flight Design.

Flight Crew (NASA/JSC)
* Issue with the Knife Edge Seals–NESC, Booster Controllers and CB got together and are satisfied with knife edge seals for STS-121 and 114. Have some long term recommendations but looks like everyone is coming together. Mr Parsons commented that Boeing had an internal Independent Assessment and had a similar finding and are ok with the first couple of missions.

System Engineering Integration (NASA/JSC)
* MSFC Redstone, the testing on ice liberation–finished up the feedline bellows testing with the final hard ice run and the drip lip configuration. Did a run with 90% humidity and had one small 2 inch piece of ice liberated. This is not a statistically significant dataset but appears that the harder the ice, the less liberates. ET and Marshall Engineering are doing good job.
* Kubrick–made first attempt at putting ice into a Mach 3.5 Wind tunnell. It is very complicated to eject ice at just the right moment to get the right data. This is complex operation to get to work.
* The MAF TIM went well and want to thank ET for hosting, NESC, ITA agreed on distributions and came up with a couple of different distributions to run on different parameters such as on voids, fracture mechanics and came to some resolution. They scheduled tagup to finalize it. Thinks that the End to End TIM may slip a little bit but does not think that the DVR will slip.
* Trying to get the I-Load owners geared up to look at implications of flying STS-121 stack on STS-114. Requests that the I-load owners recognize the fact and get potential impacts to Darrell Stamper and if any rebuilding needed to Flight Software if we have to swap stacks.

Strategic Planning (NASA/JSC)
* Hostng Headquarters folks today in a Midterm Review of Space Operations demonstrations being put together in the labs like avionics and some performance results being worked with Shuttle Derived. They will be here today.
* Next week, will be going to Canoga Park to participate in the Shuttle Derived Vehicle TIM. This will be the last TIM for Design Cycle 3 and then will go a ‘Road Show’ to each of the Centers. Will be at JSC on June 9th. Mostly for Civil Servants review.

* PRCB today at 9am Central has 6 items on the agenda…2 Integrated Hazard Reports, 3 Waivers, and a discussion of the TPS Damage Assessment Model

* The Russian trip has been cancelled. This was a trip to KSC.

USA Management (USA/JSC)
* Meeting at 10am Central to discuss launch options and will bring to 3pm meeting if anything big comes out.

At Dentist today.

In listening to the MMT Sim Debrief, thinks that we accomplished all the objectives and everything went well. Chuck Shaw commented that everyone got a lot of benefit from the sim and everyone wrestled with the problems that the RTF Task Group wanted us to do. Believes that the MMT wrestled with it correctly and
showed they are a pretty coherent team.

John Casper commented that the RTF Task Group said that they got everything they wanted from the sim and thought that it went very well.

Related Articles